# Detection of Correlated Random Vectors

### v1

@@ 1 Introduction
The above setting is motivated by a recent line of work on what is known as the data alignment problem, introduced and explored in, e.g., [CMK18, DCK19]. In this problem, $\mathsf{X}$ and $\mathsf{Y}$ are $\mathbb{R}^{n\times d}$ random matrices, with $n$ typically designating the number of users, each with $d$ features. As in the problem above, there is an unknown (or randomly generated) permutation that matches users in $\mathsf{X}$ with those in $\mathsf{Y}$. When a pair of entries from these databases is matched, their features are dependent according to a known distribution, whereas for unmatched entries, the features are independent. In the recovery problem, the goal is to recover the unknown permutation. When the two databases have independent standard normal entries, with correlation coefficient $\rho$ between the entries of matched rows, it has been demonstrated in [DCK19] that perfect recovery is attainable if $\rho^{2}=1-o(n^{-4/d})$, while it becomes impossible if $\rho^{2}=1-\omega(n^{-4/d})$ as both $n$ and $d$ tend to infinity.
@@
The detection problem, which is more relevant to our paper, has also undergone extensive investigation in [KN22a, KN22b, Tam23, EH23], in the Gaussian case, and recently, for general distributions in [PH23]. In a nutshell, in the regime where $d\to\infty$, if $\rho^{2}d\to 0$ weak detection (performing slightly better than random guessing) is information-theoretically impossible, independently of the value of $n$, while if $\rho^{2}d\to\infty$, efficient strong detection (with vanishing error probability) is possible. While the statistical limits for detection and recovery are clear when $d\to\infty$, the case where $d$ is fixed is still a mystery. Specifically, for $d\geq d_{0}$, and some $d_{0}\in\mathbb{N}$, it was shown in [EH23] that if $\rho^{2}=1-o(n^{-2/(d-1)})$ then strong detection is possible, while if $\rho\leq\rho^{\star}(d)$, for some function $\rho^{\star}(d)$ of $d$, then strong detection is impossible. In particular, for $d=1$ we have $\rho^{\star}(1)=1/2$, while an algorithmic upper bound in this case is missing. Accordingly, this evident substantial gap between the current known upper and lower bounds, even in the elementary case of $d=1$, sets the main goal of our paper.
@@
The main results of our work are sharp information-theoretic thresholds for the impossibility and possibility of strong detection in the one-dimensional case, thereby closing the aforementioned gap in [EH23, KN22b, PH23, Tam23]. Specifically, we prove that strong detection is impossible when $\rho^{2}$ is bounded away from $1$ and possible when $\rho^{2}=1-o(n^{-4})$. Another significant aspect of our results concerns with the proof technique of our lower bound, in which we abandon the standard approach of upper bounding the second moment of the likelihood ratio directly. Instead, we use an orthogonal decomposition of the likelihood ratio with respect to (w.r.t.) Hermite polynomials, and compute the second moment using Parseval’s equality. Not only that this new approach enables us to improve upon the $\rho^{\star}$ threshold from [EH23], and close the gap, it also reveals an intriguing connection between random permutations and integer partition functions.
@@
Then, in the second part of our paper, we use the techniques established for the one-dimensional, and study a generalized version, where $\mathsf{X}$ and $\mathsf{Y}$ are now $n\times d$ databases/matrices, and further consider the scenario of partial correlation. To wit, under the alternative hypothesis, only a random subset of $k$ out of the $n$ rows of $\mathsf{X}$ are correlated with some $k$ rows of a randomly permuted version of $\mathsf{Y}$. This aims to model the natural and practical scenario where in two different databases (e.g., movie rating datasets) only a subset of the total population of users in each database, is common. For this model, we prove non-trivial thresholds for impossibility and possibility of weak and strong detection w.r.t. the parameters $\rho,d,n$ and $k$. For example, in the fully correlated case where $k=n$, our results coincide with the sharp thresholds proved in [EH23]. When $k=O(\log n)$, our results suggest that detection becomes statistically harder compared to the fully correlated model, regardless of the value of $d$.
@@
We will now present a brief overview of a few related works. The problem of partially recovering the hidden alignment was investigated in [DCK20]. In [SGE19] necessary and sufficient conditions for successful recovery using a typicality-based framework were developed. Furthermore, [BE21] and [BE22] addressed alignment recovery in scenarios involving feature deletions and repetitions, respectively. Finally, the task of joint detection and recovery was proposed and analyzed in [Tam23]. It is noteworthy that the difficulties in database alignment and detection are closely linked to various planted matching problems, particularly the graph alignment problem. This problem involves detecting edge correlations between two random graphs with unlabeled nodes, see, e.g., [MMX21, DWXY21, PG11, DMWX18, WXY20, MWXY23, WXY22, Gan20].
@@ 2 Model Formulation
Preliminaries and notation.
Throughout this work, we use lower case letters $x$ for scalars, upper case letters $X$ for random variables. We denote the set of integer $\left\{1,2,\dots n\right\}$ by $[n]$, and with abuse of notation, we use $\binom{[n]}{k}$ in order to denote the set of all subsets of $[n]$ of size $k$. We denote the set of all permutation over $[n]$ by $\mathbb{S}_{n}$, and for $\sigma\in\mathbb{S}_{n}$, we let $\sigma(i)$ denote the value to which $\sigma$ maps $i\in[n]$. For two functions $f,g:\mathbb{N}\to\mathbb{R}$ we say that $f=o(g)$ if $f(n)/g(n)\xrightarrow[]{n\to\infty}0$, and $f=O(g)$ if there exists a constant $C$ such that $f(n)\leq C\cdot g(n)$ for all $n\in\mathbb{N}$. By the same convention, we say that $f=\omega(g)$ if $g=o(f)$ and $f=\Omega(g)$ if $g=O(f)$. For two measures $\mu$ and $\nu$ on the same measurable space such that $\nu$ is absolutely continuous w.r.t. $\mu$ (that is, $\mu(A)=0\implies\nu(A)=0$ for all $A$), we denote the Radon-Nikodym derivative of $\nu$ w.r.t. $\mu$ by $\frac{d\mu}{d\nu}$. We use ${\cal N}(\mathbf{\mu},\Sigma)$ to denote the normal distribution with expectation $\mathbf{\mu}$ and covariance matrix $\Sigma$. For $d\in\mathbb{N}$ we use ${\cal N}(\mathbf{\mu},\Sigma)^{\otimes d}$ to denote the independent product of $d$ ${\cal N}(\mathbf{\mu},\Sigma)$-Normal distributions.
Probabilistic model.
As described in the introduction, we deal with the following hypothesis testing problem,
@@
Problem formulation.
A test function for our problem is a function $\phi:\mathbb{R}^{n}\times\mathbb{R}^{n}\to\{0,1\}$, designed to determine which of the hypothesis ${\cal H}_{0},{\cal H}_{1}$ occurred. The risk of a test $\phi$ is defined as the sum of its average Type-I and Type-II error probabilities, i.e.,
@@
While admissibility of strong detection clearly refers to the existence of algorithms that correctly detects with probability that tends to $1$, weak detection implies the the existence of algorithms which are asymptotically better then randomly guessing which of the hypothesis occurred. The next two sections are devoted to the investigation of the one-dimensional detection problem in (1). As mentioned in the Introduction, we investigate also a generalization of (1) to high-dimensions and partial correlations (see, Section 5).
@@ 3 Phase Transition in 1D
Lower bounds.
We have the following impossibility guarantees for strong and weak detection, respectively.
@@
Consider the detection problem in (1). For any sequence $(\rho,n)=(\rho_{k},n_{k})_{k}$ such that $\rho^{2}=1-\Omega(1)$, we have
$\mathsf{R}^{\star}=1-\Omega(1).$ (7)
@@
Consider the detection problem in (1). For any sequence $(\rho,n)=(\rho_{k},n_{k})_{k}$ such that $\rho^{2}=o(1)$, we have
@@
Based on Theorem 1 we see that strong detection is impossible for any fixed $\rho^{2}<1$. This improves strictly the bound in [EH23], where it was shown that strong detection is impossible when $\rho^{2}<1/2$. Below we show that strong detection is possible if $\rho^{2}$ approaches unity sufficiently fast and that weak detection is possible when $\rho^{2}$ is bounded away from zero.
Upper bounds.
Next, we present our upper bounds, staring with a strong detection guarantee. To that end, we now define a testing procedure and analyze its performance. For $x,y\in\mathbb{R}$, let the individual likelihood of $x$ and $y$ be defined by
${\cal L}_{\mathsf{I}}(x,y)\triangleq-\frac{1}{2}\log(1-\rho^{2})-\frac{\rho^{2}}{2(1-\rho^{2})}+\frac{\rho}{1-\rho^{2}}xy,$ (9)
@@
where ${\cal I}(X_{i},Y_{j})\triangleq\mathbbm{1}\left\{{\cal L}_{\mathsf{I}}(X_{i},Y_{j})\geq\tau_{\mathsf{count}}\right\}$, ${\cal P}_{\rho}\triangleq\mathbb{P}_{\rho}\left[{\cal L}_{\mathsf{I}}(A,B)\geq\tau_{\mathsf{count}}\right]$, $\tau_{\mathsf{count}}\in\mathbb{R}$, and
$\mathbb{P}_{\rho}\triangleq\mathcal{N}\left(\begin{bmatrix}0\\ 0\end{bmatrix},\begin{bmatrix}1&\rho\\ \rho&1\end{bmatrix}\right).$ (15)
We also define ${\cal Q}_{\rho}\triangleq\mathbb{P}_{0}\left[{\cal L}_{\mathsf{I}}(A,B)\geq\tau_{\mathsf{count}}\right]$, where $\mathbb{P}_{0}\equiv\mathbb{P}_{\rho=0}$. Roughly speaking, $\phi_{\mathsf{count}}$ counts the number of pairs whose likelihood individually exceed a certain threshold. To present our main result, we need the following definitions. For simplicity of notation we denote $Q_{XY}={\cal N}(\mathbf{0},\Sigma_{0})$ and $P_{XY}={\cal N}(\mathbf{0},\Sigma_{\rho})$. For $\theta\in(-d_{\mathsf{KL}}(Q_{XY}||P_{XY}),d_{\mathsf{KL}}(P_{XY}||Q_{XY}))$, we define the Chernoff’s exponents $E_{P},E_{Q}:\mathbb{R}\to[-\infty,\infty)$ as the Legendre transforms of the log-moment generating functions, namely,
$E_{Q}(\theta)$ $\triangleq\sup_{\lambda\in\mathbb{R}}\lambda\theta-\psi_{Q}(\lambda),$ (16)
$E_{P}(\theta)$ $\triangleq\sup_{\lambda\in\mathbb{R}}\lambda\theta-\psi_{P}(\lambda),$ (17)
where $\psi_{Q}(\lambda)\triangleq\log\mathbb{E}_{Q}[\exp(\lambda{\cal L}_{\mathsf{I}})]$ and $\psi_{P}(\lambda)\triangleq\log\mathbb{E}_{P}[\exp(\lambda{\cal L}_{\mathsf{I}})]$. Then, applying standard Chernoff’s bound on these tails yield that if the threshold $\tau_{\mathsf{count}}$ is such that $\tau_{\mathsf{count}}\in(-d_{\mathsf{KL}}(Q_{XY}||P_{XY}),d_{\mathsf{KL}}(P_{XY}||Q_{XY}))$, then
@@
Consider the detection problem in (1), and the count test in (10). Suppose there is a $\tau_{\mathsf{count}}\in(-d_{\mathsf{KL}}(Q_{XY}||P_{XY}),d_{\mathsf{KL}}(P_{XY}||Q_{XY}))$ with
@@
Then, $\mathsf{R}(\phi_{\mathsf{count}})\to 0$, as $n\to\infty$. In particular, for (3) to hold it is suffice that $\rho^{2}=1-o(n^{-4})$.
Next, we consider weak detection. To that end, let $\theta\in\mathbb{R}_{+}$, and define the test,
@@
Such a value exists by the definition of the total-variation distance for centered Gaussian random variables (see, e.g., [DMR23, pg. 10]). We then have the following result.
@@
Consider the detection problem in (1), and the comparison test in (20), with $\theta$ given by . If $\rho^{2}=\Omega(1)$, then $\lim_{n\to\infty}\mathsf{R}(\phi_{\mathsf{comp}})<1$.
We provide the proofs of Theorems 3 and 4 in Subsection 5.2, where in fact, we prove a more general result for a high-dimensional model which allows for partial correlations, as described in Section 5. We would like to mention here that the upper bounds in [EH23] exclude the case where $d=1$, and as so the above result is novel. Furthermore, Theorem 4 is also novel; in particular, the weak detection results in [KN22b, EH23] again exclude the $d=1$ case (and in fact a range of values of $d$, as discussed in more detail in Section 5). Finally, we would like to discuss the statistical gap between the lower and upper bounds in Theorem 1 and Theorem 3, respectively. While according to Theorem 1 for strong detection to be possible, $\rho^{2}$ must converge to unity, as $n\to\infty$, it is not clear yet what is the sufficient rate of convergence. We suspect that Theorem 3 is not optimal, and that, slower rates are possible. An intriguing question is to analyze the optimal Neyman-Pearson test, which seems quite challenging similarly to other related problems where a latent combinatorial structure is planted.

### v3

@@ 1 Introduction
The above setting is motivated by a recent line of work on what is known as the data alignment problem, introduced and explored in, e.g., [1, 2]. In this problem, $\mathsf{X}$ and $\mathsf{Y}$ are $\mathbb{R}^{n\times d}$ random matrices, with $n$ typically designating the number of users, each with $d$ features. As in the problem above, there is an unknown (or randomly generated) permutation that matches users in $\mathsf{X}$ with those in $\mathsf{Y}$. When a pair of entries from these databases is matched, their features are dependent according to a known distribution, whereas for unmatched entries, the features are independent. In the recovery problem, the goal is to recover the unknown permutation. When the two databases have independent standard normal entries, with correlation coefficient $\rho$ between the entries of matched rows, it has been demonstrated in [2] that perfect recovery is attainable if $\rho^{2}=1-o(n^{-4/d})$, while it becomes impossible if $\rho^{2}=1-\omega(n^{-4/d})$ as both $n$ and $d$ tend to infinity.
@@
The detection problem, which is more relevant to our paper, has also undergone extensive investigation in [3, 4, 5, 6], in the Gaussian case, and recently, for general distributions in [7]. In a nutshell, in the regime where $d\to\infty$, if $\rho^{2}d\to 0$ weak detection (performing slightly better than random guessing) is information-theoretically impossible, independently of the value of $n$, while if $\rho^{2}d\to\infty$, efficient strong detection (with vanishing error probability) is possible. While the statistical limits for detection and recovery are clear when $d\to\infty$, the case where $d$ is fixed is still a mystery. Specifically, for $d\geq d_{0}$, and some $d_{0}\in\mathbb{N}$, it was shown in [6] that if $\rho^{2}=1-o(n^{-2/(d-1)})$ then strong detection is possible, while if $\rho\leq\rho^{\star}(d)$, for some function $\rho^{\star}(d)$ of $d$, then strong detection is impossible. In particular, for $d=1$ we have $\rho^{\star}(1)=1/2$, while an algorithmic upper bound in this case is missing. Accordingly, this evident substantial gap between the current known upper and lower bounds, even in the elementary case of $d=1$, sets the main goal of our paper.
@@
The main results of our work are sharp information-theoretic thresholds for the impossibility and possibility of strong detection in the one-dimensional case, thereby closing the aforementioned gap in [6, 4, 7, 5]. Specifically, we prove that strong detection is impossible when $\rho^{2}$ is bounded away from $1$ and possible when $\rho^{2}=1-o(n^{-4})$. Another significant aspect of our results concerns with the proof technique of our lower bound, in which we abandon the standard approach of upper bounding the second moment of the likelihood ratio directly. Instead, we use an orthogonal decomposition of the likelihood ratio with respect to (w.r.t.) Hermite polynomials, and compute the second moment using Parseval’s identity. Not only that this new approach enables us to improve upon the $\rho^{\star}$ threshold from [6], and close the gap, it also reveals an intriguing connection between random permutations and integer partition functions.
@@
Then, in the second part of our paper, we use the techniques established for the one-dimensional, and study a generalized version, where $\mathsf{X}$ and $\mathsf{Y}$ are now $n\times d$ databases/matrices, and further consider the scenario of partial correlation. To wit, under the alternative hypothesis, only a random subset of $k$ out of the $n$ rows of $\mathsf{X}$ are correlated with some $k$ rows of a randomly permuted version of $\mathsf{Y}$. This aims to model the natural and practical scenario where in two different databases (e.g., movie rating datasets) only a subset of the total population of users in each database, is common. For this model, we prove non-trivial thresholds for impossibility and possibility of weak and strong detection w.r.t. the parameters $\rho,d,n$ and $k$. For example, in the fully correlated case where $k=n$, our results coincide with the sharp thresholds proved in [6]. When $k=O(\log n)$, our results suggest that detection becomes statistically harder compared to the fully correlated model, regardless of the value of $d$.
@@
We will now present a brief overview of a few related works. The problem of partially recovering the hidden alignment was investigated in [8]. In [9] necessary and sufficient conditions for successful recovery using a typicality-based framework were developed. Furthermore, [10] and [11] addressed alignment recovery in scenarios involving feature deletions and repetitions, respectively. Finally, the task of joint detection and recovery was proposed and analyzed in [5]. We also mention a series of papers [12, 13, 14, 15], in which a variant the database alignment problem (i.e., recovery) was considered. Specifically, these papers deal with the task of matching random sequences to their distinct generative distributions, which are either known or unknown but a training sequence from each database is available. While this is similar in nature to the database alignment problem, the techniques and results are quite different. It is noteworthy that the difficulties in database alignment and detection are closely linked to various planted matching problems, particularly the graph alignment problem. This problem involves detecting edge correlations between two random graphs with unlabeled nodes, see, e.g., [16, 17, 18, 19, 20, 21, 22, 23].
@@ 2 Model Formulation
Preliminaries and notation. Throughout this work, we use lower case letters $x$ for scalars, upper case letters $X$ for random variables. We denote the set of integer $\left\{1,2,\dots n\right\}$ by $[n]$, and with abuse of notation, we use $\binom{[n]}{k}$ in order to denote the set of all subsets of $[n]$ of size $k$. We denote the set of all permutation over $[n]$ by $\mathbb{S}_{n}$, and for $\sigma\in\mathbb{S}_{n}$, we let $\sigma(i)$ denote the value to which $\sigma$ maps $i\in[n]$. For two functions $f,g:\mathbb{N}\to\mathbb{R}$ we say that $f=o(g)$ if $f(n)/g(n)\xrightarrow[]{n\to\infty}0$, and $f=O(g)$ if there exists a constant $C$ such that $f(n)\leq C\cdot g(n)$ for all $n\in\mathbb{N}$. By the same convention, we say that $f=\omega(g)$ if $g=o(f)$ and $f=\Omega(g)$ if $g=O(f)$. For two measures $\mu$ and $\nu$ on the same measurable space such that $\nu$ is absolutely continuous w.r.t. $\mu$ (that is, $\mu(A)=0\implies\nu(A)=0$ for all $A$), we denote the Radon-Nikodym derivative of $\nu$ w.r.t. $\mu$ by $\frac{\mathrm{d}\mu}{\mathrm{d}\nu}$. We use ${\cal N}(\mathbf{\mu},\Sigma)$ to denote the normal distribution with expectation $\mathbf{\mu}$ and covariance matrix $\Sigma$. For $d\in\mathbb{N}$ we use ${\cal N}(\mathbf{\mu},\Sigma)^{\otimes d}$ to denote the independent product of $d$ ${\cal N}(\mathbf{\mu},\Sigma)$-Normal distributions.
Probabilistic model. As described in the introduction, we deal with the following hypothesis testing problem,
@@
Problem formulation. A test function for our problem is a function $\phi:\mathbb{R}^{n}\times\mathbb{R}^{n}\to\{0,1\}$, designed to determine which of the hypothesis ${\cal H}_{0},{\cal H}_{1}$ occurred. The risk of a test $\phi$ is defined as the sum of its average Type-I and Type-II error probabilities, i.e.,
@@
While admissibility of strong detection clearly refers to the existence of algorithms that correctly detects with probability that tends to $1$, weak detection implies the the existence of algorithms which are asymptotically better then randomly guessing which of the hypothesis occurred. The next two sections are devoted to the investigation of the one-dimensional detection problem in (LABEL:eqn:decproblem). As mentioned in the Introduction, we investigate also a generalization of (LABEL:eqn:decproblem) to high-dimensions and partial correlations (see, Section 5).
@@ 3 Phase Transition in 1D
Lower bounds. We have the following impossibility guarantees for strong and weak detection, respectively.
@@
Consider the detection problem in (LABEL:eqn:decproblem). For any sequence $(\rho,n)=(\rho_{k},n_{k})_{k}$ such that $\rho^{2}=1-\Omega(1)$, we have
$\mathsf{R}^{\star}=\Omega(1).$ (5)
@@
Consider the detection problem in (LABEL:eqn:decproblem). For any sequence $(\rho,n)=(\rho_{k},n_{k})_{k}$ such that $\rho^{2}=o(1)$, we have
@@
Let us discuss the above results. From Theorem 1 we see that strong detection is statistically impossible if $\rho^{2}$ is bounded away from $1,$ and in particular, for any fixed $\rho^{2}<1$. This improves strictly the known bound in [6], where it was shown that strong detection is statistically impossible whenever $\rho^{2}<1/2$. As for weak detection, our lower bound in Theorem 2 implies that weak detection is statistically impossible whenever the correlation $\rho^{2}$ decays to zero; this matches the results in [6].
Upper bounds. Next, we present our upper bounds, starting with a strong detection guarantee. To that end, we now define a testing procedure and analyze its performance. For simplicity of notation, we denote $Q_{XY}={\cal N}(\mathbf{0},\Sigma_{0})$ and $P_{XY}={\cal N}(\mathbf{0},\Sigma_{\rho})$, and the associated joint probability density functions by $f_{P}$ and $f_{Q}$, respectively. Then, for $x,y\in\mathbb{R}$, let the individual likelihood of $x$ and $y$ be defined by
${\cal L}_{\mathsf{I}}$ $(x,y)\triangleq\log\frac{f_{P}(x,y)}{f_{Q}(x,y)}$ (7)
$=-\frac{1}{2}\log(1-\rho^{2})-\frac{\rho^{2}}{2(1-\rho^{2})}(x^{2}+y^{2})+\frac{\rho}{1-\rho^{2}}xy,$ (8)
@@
where ${\cal I}(X_{i},Y_{j})\triangleq\mathbbm{1}\left\{{\cal L}_{\mathsf{I}}(X_{i},Y_{j})\geq\tau_{\mathsf{count}}\right\}$, ${\cal P}_{\rho}\triangleq\mathbb{P}_{P}\left[{\cal L}_{\mathsf{I}}(A,B)\geq\tau_{\mathsf{count}}\right]$, and $\tau_{\mathsf{count}}\in\mathbb{R}$, where $\mathbb{P}_{P}(\mathcal{F})$ means that the probability of the event $\mathcal{F}$ is evaluated w.r.t. the joint distribution $P_{XY}$. Similarly, we also define ${\cal Q}_{\rho}\triangleq\mathbb{P}_{Q}\left[{\cal L}_{\mathsf{I}}(A,B)\geq\tau_{\mathsf{count}}\right]$, where the probability is evaluated w.r.t. the joint distribution $Q_{XY}$. Roughly speaking, $\phi_{\mathsf{count}}$ counts the number of pairs whose likelihood individually exceed the threshold $\frac{1}{2}n{\cal P}_{\rho}$. Intuitively, since the individual likelihoods are proportional to the products of the entries of $\mathsf{X}$ and $\mathsf{Y}$, we expect the count to be larger under the alternative hypothesis, when these vectors are correlated. While the choice of the threshold $\frac{1}{2}n{\cal P}_{\rho}$ is not unique (and perhaps even not optimal), it allows us to control the Type-I and Type-II error probabilities. Indeed, intuitively, when $\rho^{2}$ is sufficiently close to unity (which as will be seen below is the relevant regime where the count test is successful) we get that ${\cal Q}_{\rho}=o(n^{2}{\cal P}_{\rho})$, and therefore, the summation of ${\cal L}_{\mathsf{I}}(X_{i},Y_{i})$ over uncorrelated pairs $\{(X_{i},Y_{j})\}_{i,j}$ is “relatively small” with high probability (e.g., by the law of large numbers); this eliminates errors of Type-I. Moreover, we have that ${\cal P}_{\rho}$ is strictly positive, and thus the summation of ${\cal L}_{\mathsf{I}}(X_{i},Y_{i})$ over correlated pairs $\{(X_{i},Y_{j})\}_{i,j}$ is strictly positive (again, by the law of large numbers), which eliminates errors of Type-II. It should be mentioned here that a similar test was proposed in [6]. Specifically, in [6], the number of normalized inner products is being counted, while here we count the inner products directly. As it turns out, the normalization procedure in [6] excludes the $d=1$ case we consider in this section.
To present our main result, we need the following definitions. For $\theta\in(-d_{\mathsf{KL}}(Q_{XY}||P_{XY}),d_{\mathsf{KL}}(P_{XY}||Q_{XY}))$, we define the Chernoff’s exponents $E_{P},E_{Q}:\mathbb{R}\to[-\infty,\infty)$ as the Legendre transforms of the log-moment generating functions, namely,
$E_{Q}(\theta)$ $\triangleq\sup_{\lambda\in\mathbb{R}}\left[\lambda\theta-\psi_{Q}(\lambda)\right],$ (10)
$E_{P}(\theta)$ $\triangleq\sup_{\lambda\in\mathbb{R}}\left[\lambda\theta-\psi_{P}(\lambda)\right],$ (11)
where $\psi_{Q}(\lambda)\triangleq\log\mathbb{E}_{Q}[\exp(\lambda{\cal L}_{\mathsf{I}}(A,B))]$ and $\psi_{P}(\lambda)\triangleq\log\mathbb{E}_{P}[\exp(\lambda{\cal L}_{\mathsf{I}}(A,B))]$. Closed-form expressions for $\psi_{Q}$ and $\psi_{P}$ in our setting, can be found in Lemma 5. Then, applying standard Chernoff’s bound on these tails yield that if the threshold $\tau_{\mathsf{count}}$ is such that $\tau_{\mathsf{count}}\in(-d_{\mathsf{KL}}(Q_{XY}||P_{XY}),d_{\mathsf{KL}}(P_{XY}||Q_{XY}))$, then
@@
Consider the detection problem in (LABEL:eqn:decproblem), and the count test in (9). Suppose there is a $\tau_{\mathsf{count}}\in(-d_{\mathsf{KL}}(Q_{XY}||P_{XY}),d_{\mathsf{KL}}(P_{XY}||Q_{XY}))$ with
@@
Then, $\mathsf{R}(\phi_{\mathsf{count}})\to 0$, as $n\to\infty$. In particular, for (13) to hold it suffices that $\rho^{2}=1-o(n^{-4})$.
Recall that in Theorem 1 we proved that strong detection is statistically impossible for any $\rho^{2}<1$; therefore, for strong detection to be possible $\rho^{2}$ must converge to unity. Theorem 3 shows that indeed the count test that achieves strong detection if $\rho^{2}$ approaches unity sufficiently fast. We would like to mention here that there is still a statistical gap between the lower and upper bounds in Theorem 1 and Theorem 3, respectively. While according to Theorem 1 for strong detection to be possible, $\rho^{2}$ must converge to unity, as $n\to\infty$, it is not clear yet what is the sufficient rate of convergence. We suspect that Theorem 3 is not optimal, and that, slower rates are possible. An intriguing question is to analyze the optimal Neyman-Pearson test, which seems quite challenging similarly to other related problems where a latent combinatorial structure is planted. Finally, we would like to mention here that the strong detection upper bounds in [4, 6] exclude the case where $d=1$, and as so Theorem 3 is novel.
Next, we consider weak detection. By definition, any test that achieves strong detection achieves weak detection automatically. As so, the count test above achieves weak detection under the same conditions stated in Theorem 3. In fact, it is rather straightforward to show that weak detection is possible using the count test even if, for example, the condition $\rho^{2}=1-o(n^{-4})$ is replaced by the weaker condition $\rho^{2}=1-\mathsf{C}\cdot n^{-4}$, for some constant $\mathsf{C}>0$ (or, more generally, the $\omega$-asymptotic in (13) is replaced by some constant $\mathsf{C}>0$). Nonetheless, for weak detection, it turns out that the following rather simpler test exhibits better performance guarantees. Let $\theta\in\mathbb{R}_{+}$, and define the test,
@@
Such a value exists by the definition of the total-variation distance for centered Gaussian random variables (see, e.g., [24, pg. 10]). The intuition behind this choice of $\theta$ is that, as it turns out, the right-hand-side (r.h.s.) of (15) is exactly the reward (i.e., $1-\mathsf{R}(\phi)$) associated with the comparison test. Thus, since the total-variation distance at the left-hand-side of (15) is clearly positive for any $\rho^{2}=\Omega(1)$, we obtain that the risk is strictly bounded by unity. We then have the following result.
@@
Consider the detection problem in (LABEL:eqn:decproblem), and the comparison test in (14), with $\theta$ given by (15). If $\rho^{2}=\Omega(1)$, then $\lim_{n\to\infty}\mathsf{R}(\phi_{\mathsf{comp}})<1$.
Comparing Theorems 3 and 4 we see that for weak detection we can construct a test for which it is suffice that $\rho^{2}$ is of order constant (i.e., bounded away from zero), while for strong detection our test requires the correlation to converge to unity sufficiently fast. We would like to emphasize that Theorem 4 is novel; in particular, the weak detection results in [4, 6] again exclude the $d=1$ case (and in fact a range of values of $d$, as discussed in more detail in Section 5). We provide the proofs of Theorems 3 and 4 in Section 5.2, where in fact, we prove a more general result for a high-dimensional model which allows for partial correlations, as described in Section 5. We summarize the above results in Table 1, next to previously known bounds in the literature. Finally, we conclude this subsection by showing a numerical evaluation of our count test. Specifically, in Fig. 2, we present the empirical risk, averaged over $10^{3}$ Monte-Carlo runs, associated with the count test in (9) with $\tau_{\mathsf{count}}=0$, as a function of $\rho$, for $n=100$. As predicted by our theoretical results, it can be seen that if the correlation is sufficiently close to unity, then the associated risk decreases.
Weak Detection Strong Detection
Current/previous work Possible Impossible Possible Impossible
[…]
Previous work $o(1)$- $\frac{1}{2}$
Table 1: A summary of our bounds on $\rho^{2}$, for weak and strong detection in the one-dimensional case, compared with the results of [6].
[…]
Figure 2: The risk of the count test $\phi_{\mathsf{count}}$ as a function of $\rho$, for $d=1$, $n=100$ and values of $\rho^{2}$ spanning between $1-n^{-2}$ and $1-n^{-5}$. The results are compatible with Theorem 1, showing that the risk indeed vanishes when $\rho^{2}=1-o(n^{-4}$).